# Hierarchical Uncertainty-Aware Graph Neural Network

### v1

@@ 7. Conclusion
This paper investigates how hierarchical representations and uncertainty estimation jointly affect the robustness of Graph Neural Networks (GNNs). We analyze uncertainty propagation at three structural scales: nodes, communities, and a global context. Also, we prove that the resulting message-passing operator enjoys tighter PAC-Bayes generalization bounds and a contractive update that guarantees convergence on arbitrary graphs. The proposed method adaptively re-weights messages with scale-specific uncertainty scores and dynamically shifts attention from unreliable one-hop neighbors to more trustworthy community or global evidence. Extensive experiments under random noise and adaptive attacks show that HU-GNN loses only a small portion of accuracy on average, delivering state-of-the-art results on both homophilic and heterophilic benchmarks. These findings demonstrate that coupling a multi-level structure with learned uncertainty is a powerful remedy for over-smoothing and adversarial fragility.

### v2

@@ 7. Conclusion
This paper investigates how hierarchical representations and uncertainty estimation jointly affect the robustness of GNNs. We analyze uncertainty propagation at three structural scales: nodes, communities, and a global context. Also, we prove that the resulting message-passing operator enjoys tighter PAC-Bayes generalization bounds and a contractive update that guarantees convergence on arbitrary graphs. The proposed method adaptively re-weights messages with scale-specific uncertainty scores and dynamically shifts attention from unreliable one-hop neighbors to more trustworthy community or global evidence. Extensive experiments under random noise and adaptive attacks show that HU-GNN loses only a small portion of accuracy on average, delivering state-of-the-art results on both homophilic and heterophilic benchmarks. The findings demonstrate that coupling a multi-level structure with learned uncertainty is a powerful remedy for over-smoothing and adversarial fragility.